OpenStudy (anonymous):

factor by grouping. 2+2y+x+xy

OpenStudy (raden):

2+2y+x+xy = 2(1+y) + x(1+y) = (1+y)(2+x)

OpenStudy (anonymous):

Thanks(:

OpenStudy (raden):

welcome